OpenStudy (yessenia27):

Evan spent 20 hours doing homework last week. This week he spent 25 hours doing homework. He says that he spent 125% more time doing homework this week. Is he correct? Show your work to justify your decision.

MARC (marc_d):

\[\frac{ this~week }{ last~week }\times100percent\]

MARC (marc_d):

\[\frac{ 25 }{ 20 }\times100percent\]

MARC (marc_d):

\[=\frac{ 5 }{ 4 }\times100percent\]

MARC (marc_d):

\[=125percent\]

MARC (marc_d):

Therefore,he is correct.

MARC (marc_d):

@yessenia27